# Quantum amplification and simulation of strong and ultrastrong coupling of light and matter

### v1

@@ 3.1.3 Implementations in trapped thermal atoms
An implementation that followed the original proposal of Ref. Dimer et al. (2007) more closely, by using the atomic electronic states instead of the atomic motional states to encode a two-level system, was shown for the first time in Ref. Baden et al. (2014), where the effective two-level system was encoded by the hyperfine ground states $\left|{F=1,m_{F}=1}\right\rangle$ and $\left|{F=2,m_{F}=2}\right\rangle$ of ${}^{87}$b atoms, with a Zeeman splitting (on the order of MHz) induced by an external magnetic field. In Ref. Zhiqiang et al. (2017), the same group further explored the rich phase diagram emergent under situations of the asymmetric coupling, i.e., $\lambda_{s}\neq\lambda_{r}$, and in Ref. Zhang et al. (2018c), the differences between co- and counter-propagating structures were discussed in detail.
The cavity-assisted Raman transitions have been investigated for a plethora of effects and their applications, such as the formation of spinor self-ordering and spin textures Kroeze et al. (2018); Landini et al. (2018), the generation of spin squeezing Dalla Torre et al. (2013) and steady-state entanglement in solid-state systems González-Tudela and Porras (2013), the simulation of supersymmetry field theories in quantum optics Tomka et al. (2015), the driven dissipative dynamics under strong symmetries (dissipative freezing) Sánchez Muñoz et al. (2019), and the engineering of spin-spin interactions mediated by photons Masson et al. (2017); Davis et al. (2019); Mivehvar et al. (2019); Norcia et al. (2018).
@@ 3.2.1 Theoretical proposals
The simulation scheme proposed in Ref. Ballester et al. (2012) considers a system consisting of a qubit, of frequency $\omega_{q}$, coupled with a strength $g$ to a cavity mode of frequency $\omega_{\rm cav}$, such that a full description of the system is given by the quantum Rabi Hamiltonian,
@@
When $g\ll\{\omega_{q},\omega_{\rm cav}\}$, the RWA can be applied to yield the simplified model described by the JC Hamiltonian
@@
It is important to note that the effective model in Eq. (130) is obtained in a rotating frame. Experimentally, some quantities in this model, like the qubit population, would thus display rapid oscillations, which could make them hard to observe. To remedy this, Ref. Ballester et al. (2012) proposes that after the two drives have been on for a time $t$, they are turned off non-adiabatically, then the qubit and the drive at $\omega_{1}$ are both detuned by some amount before the drive at $\omega_{1}$ is turned on again for the same time $t$, but with an opposite phase to before. This protocol allows one to recover the dynamics of the model in a rotating frame.
@@
Single ions can be trapped using radio-frequency fields and have some of their electronic degrees of freedom (a qubit) couple to their motion in the trap (a harmonic oscillator) through laser driving. The proposal in Ref. Pedernales et al. (2015) considers the situation when two laser drives are applied, in which case the Hamiltonian for such a setup can be written as
@@
$H=\frac{1}{2}\left(\delta_{b}-\delta_{r}\right)a^{\dagger}a+\frac{1}{4}\left(\delta_{r}+\delta_{b}\right)\sigma_{z}+\frac{i\eta\Omega}{2}\left(\sigma_{+}-\sigma_{-}\right)\left(a+a^{\dagger}\right),$ (134)
which has the form of the quantum Rabi Hamiltonian. Here, we can identify an effective harmonic-mode frequency $\left(\delta_{b}-\delta_{r}\right)/2$, an effective qubit frequency $\left(\delta_{r}+\delta_{b}\right)/2$, and an effective coupling strength $\eta\Omega/2$. These effective parameters can all be tuned by adjusting the drive frequencies and strengths, allowing the simulator to explore a wide range of parameters in the USC regime. As is pointed out in Ref. Pedernales et al. (2015), all transformations performed here commute with observables of interest (e.g., the number of excitations in the qubit and the harmonic mode), such that these observables can be directly measured in experiments. A later theory work Puebla et al. (2017) suggests that some additional laser drives may need to be added to explore the particular parameter regime where a quantum phase transition can occur.
The setup of a bichromatically driven ion can be taken further to realize other versions of the quantum Rabi model. In Ref. Felicetti et al. (2015), it was shown that it also can simulate the two-photon Rabi model, i.e., by replacing $a$ and $a^{\dagger}$ in the interaction in Eq. (126) with $a^{2}$ and $\left(a^{\dagger}\right)^{2}$, respectively, such that the photonic population only can be changed in steps of two. The key to realizing this simulation is to tune the two drives to the second sidebands instead of the first ones, i.e., to set the drive frequencies to $\omega_{b/r}=\omega_{q}\pm 2\omega_{\rm mot}+\delta_{b/r}$. Using the same assumptions as when going from Eq. (132) to Eq. (133), with these drive frequencies Eq. (132) simplifies to
@@
Further generalizations of the scheme in Ref. Pedernales et al. (2015) were proposed in Ref. Aedo and Lamata (2018). There, it was shown that the scheme can be extended to the Dicke model (the generalization of the quantum Rabi model to multiple qubits), as well as variations on that, like the biased Dicke model (which includes a term proportional to $\sigma_{x}$ for each qubit) and the anisotropic Dicke model (where the rotating and counter-rotating parts of the interaction have different strengths).
@@ 3.2.2 Experimental implementations
The first implementation of the proposal in Ref. Ballester et al. (2012) came in 2017 Braumüller et al. (2017), using superconducting circuits Gu et al. (2017); Krantz et al. (2019); Blais et al. (2021). The experimental setup, shown in Fig. 12, consisted of an LC oscillator (the harmonic mode) and a transmon qubit Koch et al. (2007) (an anharmonic LC oscillator, made anharmonic by the nonlinear inductance of Josephson junctions). Using the notation from Eq. (127), this setup had a harmonic-mode frequency $\omega_{\rm cav}/2\pi=5.948\,\mathrm{GHz}$, a qubit frequency $\omega_{q}$ tuned close to $\omega_{\rm cav}$, and a bare coupling strength $g/2\pi=4.3\,\mathrm{MHz}$, i.e., $g/\omega_{\rm cav}\lesssim 0.001$.
Figure 12: Experimental setup used in Ref. Braumüller et al. (2017) for simulation of USC with superconducting circuits. (a) Optical micrograph of the chip used, with the transmon qubit highlighted in red and the $\lambda/2$ microstrip resonator constituting the harmonic mode highlighted in blue. The resonator next to the harmonic mode is not used in the experiment; it is detuned from the other components in frequency. The scale bar in the bottom left corner is 1 mm. (b) Circuit diagram of the setup on the chip. The qubit and the resonator are capacitively coupled, and the qubit also couples capacitively to another resonator, which is probed through a transmission line (TL) to read out the qubit state. (c) Pulse sequence used to simulate the quantum Rabi Hamiltonian. (d) Illustration of the frequencies for the harmonic mode and the applied drives. Figures, with some modifications, from Ref. Braumüller et al. (2017), J. Braumüller et al., Nat. Commun. 8, 779 (2017), with permission.
Applying two drives to the qubits as explained in Sec. 3.2.1 and illustrated in Fig. 12(c,d), the experimental simulation could reach deep into the USC regime with a ratio $g_{\rm eff}/\omega_{\rm eff}\approx 0.6$ between the effective coupling strength $g_{\rm eff}$ and the effective resonance frequency $\omega_{\rm eff}$. In these conditions, the quantum revival characteristics of the USC regime were observed (they are due to a different mechanism, compared to the revivals that can occur in the JC model Eberly et al. (1980)).
@@
The proposal of Ref. Pedernales et al. (2015), discussed in Sec. 3.2.1, was first implemented in an experiment with a trapped ion in 2018 Lv et al. (2018). In that experiment, an ${}^{171}\text{Yb}^{+}$ ion was trapped and cooled down to its motional ground state. Using the notation from Eq. (131), the qubit transition frequency was $\omega_{q}/2\pi=12.642812\,\mathrm{GHz}$ (using a transition in the $S_{1/2}$ hyperfine manifold) and the harmonic motional mode had a resonance frequency of $\omega_{\rm mot}/2\pi=2.498\,\mathrm{MHz}$. The drive strength $\Omega$ of the bichromatic laser drive was chosen such that the effective coupling strength was held fixed at $g_{\rm eff}=\eta\Omega/2=2\pi\times 12.5\,\mathrm{MHz}$ and the detuning from the red sideband was set to zero, i.e., $\delta_{r}=0$. The effective resonance frequency of both the qubit and the harmonic mode then became $\omega_{\rm eff}=\delta_{b}/2$, and the detuning $\delta_{b}$ from the blue sideband was varied to achieve different ratios $g_{\rm eff}/\omega_{\rm eff}$ between 0.04 and 1.2, with the latter simulating the DSC regime.
@@
In 2021, another experimental group used the same setup with a trapped ${}^{171}\text{Yb}^{+}$ ion to study a quantum phase transition in the quantum Rabi model Cai et al. (2021), demonstrating a theoretical prediction from Ref. Hwang et al. (2015). By making the red and blue sideband detunings $\delta_{r}$ and $\delta_{b}$ almost the same, the regime with $\omega_{q}\gg\omega_{\rm mot}$ was simulated. Here, the drive strength $\Omega$ was then swept to simulate sweeping $g_{\rm eff}$ across a quantum critical point. The behaviors predicted from the quantum phase transition for both qubit and phonon populations were observed, with both populations increasing significantly when $g_{\rm eff}$ exceeded the critical value.
The next year, the same group took this setup further by simulating the Rabi–Hubbard model Mei et al. (2022). In that model, several sites, each governed by a quantum Rabi Hamiltonian, are coupled together through hopping between the bosonic modes of the respective sites. The quantum Rabi Hamiltonian was simulated as above, while the hopping interaction was given by the Coloumb interaction between ions. In the experiment, which used up to 16 ions, a quantum phase transition was observed and various dynamics of the system were studied at a scale beyond what a classical computer can simulate.
@@ 3.3 Simulated ultrastrong interactions in single-mode-driven Jaynes–Cummings systems
One of the most important hallmarks of the ultrastrong regime of light-matter coupling is the existence of nonlinear processes that do not conserve the total number of excitations in the system. In Ref. Kockum et al. (2017b), a complete description of many such processes is provided, which includes, for instance, the conversion of a single photon into multiple atomic excitations Garziano et al. (2016); Macrì et al. (2022); Tomonaga et al. (2023) or multiple photons Wang et al. (2024), multiphoton oscillations Garziano et al. (2015), converting an atomic superposition state into an entangled photonic state Macrì et al. (2018), frequency conversion Kockum et al. (2017a), and nonlinear photon-mediated interactions Stassi et al. (2017).
In order to implement the majority of these nonlinear processes, a crucial requirement is that the Hamiltonian should neither conserve the total number of excitations nor its parity. The standard quantum Rabi model, with a light-matter coupling term of the form $g(a+a^{\dagger})(\sigma^{-}+\sigma^{+})$, does not conserve the total number of excitations due to the terms $ga\sigma^{-}$ and $ga^{\dagger}\sigma^{+}$. However, these terms still keep the total parity constant, since they add or remove excitations by an even number. Therefore, most of the aforementioned processes, such as the simultaneous absorption of a single photon by several atoms Garziano et al. (2016); Tomonaga et al. (2023), require the implementation of a generalized Rabi model with the addition of another coupling term of the form $\sigma_{z}(a+a^{\dagger})$, which allows to increase or decrease excitations by an odd number. As a result, one can engineer non-linear processes characterized by the coherent oscillation between an initial state $\left|{i}\right\rangle$ and a final state $\left|{f}\right\rangle$ with a different total number of excitations, such as the simultaneous absorption of a single photon by two atoms, that would be represented by the two states $\left|{i}\right\rangle=\left|{1,g,g}\right\rangle$ and $\left|{f}\right\rangle=\left|{0,e,e}\right\rangle$. This can be achieved if the system parameters are tuned such that $\left|{i}\right\rangle$ and $\left|{f}\right\rangle$ are eigenstates of the bare Hamiltonian with the same energy, allowing one to obtain an effective Hamiltonian that generates the coherent conversion between the two states,
@@
From the point of view of simulations, a recent proposal Sánchez Muñoz et al. (2020) showed that all the nonlinear processes emerging from the generalized quantum Rabi model — including those breaking parity conservation — can be simulated in JC systems well below the USC regime, using only a single coherent drive. This result is achieved by working in the dressed basis of the atoms driven by the classical field. Let us consider the Hamiltonian of a single driven atom coupled to a cavity mode in the weak-driving regime described by the JC model, and written in the rotating frame of the drive:

### v2

@@ 3.1.3 Implementations in trapped thermal atoms
An implementation that followed the original proposal of Ref. [370] more closely, by using the atomic electronic states instead of the atomic motional states to encode a two-level system, was shown for the first time in Ref. [382], where the effective two-level system was encoded by the hyperfine ground states $\left|{F=1,m_{F}=1}\right\rangle$ and $\left|{F=2,m_{F}=2}\right\rangle$ of 87Rb atoms, with a Zeeman splitting (on the order of MHz) induced by an external magnetic field. In Ref. [383], the same group further explored the rich phase diagram emergent under situations of the asymmetric coupling, i.e., $\lambda_{s}\neq\lambda_{r}$, and in Ref. [384], the differences between co- and counter-propagating structures were discussed in detail.
The cavity-assisted Raman transitions have been investigated for a plethora of effects and their applications, such as the formation of spinor self-ordering and spin textures [385, 386], the generation of spin squeezing [319] and steady-state entanglement in solid-state systems [387], the simulation of supersymmetry field theories in quantum optics [181], the driven dissipative dynamics under strong symmetries (dissipative freezing) [388], and the engineering of spin-spin interactions mediated by photons [389, 390, 391, 392].
@@ 3.2.1 Theoretical proposals
The simulation scheme proposed in Ref. [182] considers a system consisting of a qubit, of frequency $\omega_{q}$, coupled with a strength $g$ to a cavity mode of frequency $\omega_{\rm cav}$, such that a full description of the system is given by the quantum Rabi Hamiltonian,
@@
This Hamiltonian can be viewed as an equivalent variant of the quantum Rabi Hamiltonian in the dipole gauge, i.e.,
$H_{\rm Rabi}^{\rm dg}=\omega_{\rm cav}a^{\dagger}a+\frac{\omega_{q}}{2}\sigma_{z}+ig_{\rm dg}(a^{\dagger}-a)\sigma_{x},$ (127)
since the former can be directly obtained by applying a replacement $a\rightarrow ia$ to the latter. Here, $g_{\rm dg}$ refers to the light-matter coupling strength in the dipole gauge. Note that the form of the quantum Rabi Hamiltonian in the USC regime strongly depends on the choice of gauge. In the Coulomb gauge, the standard quantum Rabi Hamiltonian is given by
$\mathcal{H}_{\rm Rabi}^{\rm cg}=\omega_{\rm cav}a^{\dagger}a+\frac{\omega_{q}}{2}\sigma_{z}+g_{\rm cg}(a+a^{\dagger})\sigma_{y}+D(a+a^{\dagger})^{2},$ (128)
where $g_{\rm cl}=g_{\rm dg}\omega_{q}/\omega_{\rm cav}$, and $D$ is the diamagnetic amplitude. According to Eq. (128), it seems as though Eq. (126) was also a simpler variant in the Coulomb gauge, with a rotation $\sigma_{y}\rightarrow\sigma_{x}$ and at the same time neglecting the diamagnetic term $\propto(a+a^{\dagger})^{2}$. However, in fact, such a system Hamiltonian in the Coulomb gauge in Eq. (128) is wrong in the USC regime, since it does not produce the correct eigenvalues and, as a result, causes the breakdown of gauge invariance [73, 74, 393]. This gauge ambiguity has been well resolved in Ref. [74] by applying the minimal-replacement rule not only to the particle kinetic energy but also to the particle nonlocal potential. In this case, the resulting system Hamiltonian in the Coulomb gauge, i.e.,
$H_{\rm Rabi}^{\rm cg}=\omega_{\rm cav}a^{\dagger}a+\frac{\omega_{q}}{2}\left\{\sigma_{z}\cos\left[2\eta(a+a^{\dagger})\right]+\sigma_{y}\sin\left[2\eta(a+a^{\dagger})\right]\right\},$ (129)
where $\eta=g_{\rm dg}/\omega_{\rm cav}$, is gauge invariant and can produce the same eigenvalues as the Hamiltonian in the dipole gauge given in Eq. (127).
When $g\ll\{\omega_{q},\omega_{\rm cav}\}$, the RWA can be applied to Eq. (126) to yield the simplified model described by the JC Hamiltonian
@@
It is important to note that the effective model in Eq. (133) is obtained in a rotating frame. Experimentally, some quantities in this model, like the qubit population, would thus display rapid oscillations, which could make them hard to observe. To remedy this, Ref. [182] proposes that after the two drives have been on for a time $t$, they are turned off non-adiabatically, then the qubit and the drive at $\omega_{1}$ are both detuned by some amount before the drive at $\omega_{1}$ is turned on again for the same time $t$, but with an opposite phase to before. This protocol allows one to recover the dynamics of the model in a rotating frame.
@@
Single ions can be trapped using radio-frequency fields and have some of their electronic degrees of freedom (a qubit) couple to their motion in the trap (a harmonic oscillator) through laser driving. The proposal in Ref. [394] considers the situation when two laser drives are applied, in which case the Hamiltonian for such a setup can be written as
@@
$H=\frac{1}{2}\left(\delta_{b}-\delta_{r}\right)a^{\dagger}a+\frac{1}{4}\left(\delta_{r}+\delta_{b}\right)\sigma_{z}+\frac{i\eta\Omega}{2}\left(\sigma_{+}-\sigma_{-}\right)\left(a+a^{\dagger}\right).$ (137)
The coupling form of $i\left(\sigma_{+}-\sigma_{-}\right)\left(a+a^{\dagger}\right)$ is an equivalent version of the light-matter coupling form of the quantum Rabi Hamiltonian in the dipole gauge [see Eq. (127)]. Here, we can identify an effective harmonic-mode frequency $\left(\delta_{b}-\delta_{r}\right)/2$, an effective qubit frequency $\left(\delta_{r}+\delta_{b}\right)/2$, and an effective coupling strength $\eta\Omega/2$. These effective parameters can all be tuned by adjusting the drive frequencies and strengths, allowing the simulator to explore a wide range of parameters in the USC regime. As is pointed out in Ref. [394], all transformations performed here commute with observables of interest (e.g., the number of excitations in the qubit and the harmonic mode), such that these observables can be directly measured in experiments. A later theory work [395] suggests that some additional laser drives may need to be added to explore the particular parameter regime where a quantum phase transition can occur.
The setup of a bichromatically driven ion can be taken further to realize other versions of the quantum Rabi model. In Ref. [396], it was shown that it also can simulate the two-photon Rabi model, i.e., by replacing $a$ and $a^{\dagger}$ in the interaction in Eq. (126) with $a^{2}$ and $\left(a^{\dagger}\right)^{2}$, respectively, such that the photonic population only can be changed in steps of two. The key to realizing this simulation is to tune the two drives to the second sidebands instead of the first ones, i.e., to set the drive frequencies to $\omega_{b/r}=\omega_{q}\pm 2\omega_{\rm mot}+\delta_{b/r}$. Using the same assumptions as when going from Eq. (135) to Eq. (136), with these drive frequencies Eq. (135) simplifies to
@@
Further generalizations of the scheme in Ref. [394] were proposed in Ref. [397]. There, it was shown that the scheme can be extended to the Dicke model (the generalization of the quantum Rabi model to multiple qubits), as well as variations on that, like the biased Dicke model (which includes a term proportional to $\sigma_{x}$ for each qubit) and the anisotropic Dicke model (where the rotating and counter-rotating parts of the interaction have different strengths).
@@ 3.2.2 Experimental implementations
The first implementation of the proposal in Ref. [182] came in 2017 [184], using superconducting circuits [10, 11, 14]. The experimental setup, shown in Fig. 12, consisted of an LC oscillator (the harmonic mode) and a transmon qubit [398] (an anharmonic LC oscillator, made anharmonic by the nonlinear inductance of Josephson junctions). Using the notation from Eq. (130), this setup had a harmonic-mode frequency $\omega_{\rm cav}/2\pi=5.948\,\mathrm{GHz}$, a qubit frequency $\omega_{q}$ tuned close to $\omega_{\rm cav}$, and a bare coupling strength $g/2\pi=4.3\,\mathrm{MHz}$, i.e., $g/\omega_{\rm cav}\lesssim 0.001$.
Figure 12: Experimental setup used in Ref. [184] for simulation of USC with superconducting circuits. (a) Optical micrograph of the chip used, with the transmon qubit highlighted in red and the $\lambda/2$ microstrip resonator constituting the harmonic mode highlighted in blue. The resonator next to the harmonic mode is not used in the experiment; it is detuned from the other components in frequency. The scale bar in the bottom left corner is 1 mm. (b) Circuit diagram of the setup on the chip. The qubit and the resonator are capacitively coupled, and the qubit also couples capacitively to another resonator, which is probed through a transmission line (TL) to read out the qubit state. (c) Pulse sequence used to simulate the quantum Rabi Hamiltonian. (d) Illustration of the frequencies for the harmonic mode and the applied drives. Figures, with some modifications, from Ref. [184], J. Braumüller et al., Nat. Commun. 8, 779 (2017), with permission.
Applying two drives to the qubits as explained in Sec. 3.2.1 and illustrated in Fig. 12(c,d), the experimental simulation could reach deep into the USC regime with a ratio $g_{\rm eff}/\omega_{\rm eff}\approx 0.6$ between the effective coupling strength $g_{\rm eff}$ and the effective resonance frequency $\omega_{\rm eff}$. In these conditions, the quantum revival characteristics of the USC regime were observed (they are due to a different mechanism, compared to the revivals that can occur in the JC model [399]).
@@
The proposal of Ref. [394], discussed in Sec. 3.2.1, was first implemented in an experiment with a trapped ion in 2018 [183]. In that experiment, an ${}^{171}\text{Yb}^{+}$ ion was trapped and cooled down to its motional ground state. Using the notation from Eq. (134), the qubit transition frequency was $\omega_{q}/2\pi=12.642812\,\mathrm{GHz}$ (using a transition in the $S_{1/2}$ hyperfine manifold) and the harmonic motional mode had a resonance frequency of $\omega_{\rm mot}/2\pi=2.498\,\mathrm{MHz}$. The drive strength $\Omega$ of the bichromatic laser drive was chosen such that the effective coupling strength was held fixed at $g_{\rm eff}=\eta\Omega/2=2\pi\times 12.5\,\mathrm{MHz}$ and the detuning from the red sideband was set to zero, i.e., $\delta_{r}=0$. The effective resonance frequency of both the qubit and the harmonic mode then became $\omega_{\rm eff}=\delta_{b}/2$, and the detuning $\delta_{b}$ from the blue sideband was varied to achieve different ratios $g_{\rm eff}/\omega_{\rm eff}$ between 0.04 and 1.2, with the latter simulating the DSC regime.
@@
In 2021, another experimental group used the same setup with a trapped ${}^{171}\text{Yb}^{+}$ ion to study a quantum phase transition in the quantum Rabi model [400], demonstrating a theoretical prediction from Ref. [401]. By making the red and blue sideband detunings $\delta_{r}$ and $\delta_{b}$ almost the same, the regime with $\omega_{q}\gg\omega_{\rm mot}$ was simulated. Here, the drive strength $\Omega$ was then swept to simulate sweeping $g_{\rm eff}$ across a quantum critical point. The behaviors predicted from the quantum phase transition for both qubit and phonon populations were observed, with both populations increasing significantly when $g_{\rm eff}$ exceeded the critical value.
The next year, the same group took this setup further by simulating the Rabi–Hubbard model [402]. In that model, several sites, each governed by a quantum Rabi Hamiltonian, are coupled together through hopping between the bosonic modes of the respective sites. The quantum Rabi Hamiltonian was simulated as above, while the hopping interaction was given by the Coulomb interaction between ions. In the experiment, which used up to 16 ions, a quantum phase transition was observed and various dynamics of the system were studied at a scale beyond what a classical computer can simulate.
@@ 3.3 Simulated ultrastrong interactions in single-mode-driven Jaynes–Cummings systems
One of the most important hallmarks of the ultrastrong regime of light-matter coupling is the existence of nonlinear processes that do not conserve the total number of excitations in the system. In Ref. [163], a complete description of many such processes is provided, which includes, for instance, the conversion of a single photon into multiple atomic excitations [171, 67, 174] or multiple photons [165], multiphoton oscillations [166], converting an atomic superposition state into an entangled photonic state [170], frequency conversion [162], and nonlinear photon-mediated interactions [85].
In order to implement the majority of these nonlinear processes, a crucial requirement is that the Hamiltonian should neither conserve the total number of excitations nor its parity. The quantum Rabi model, with a light-matter coupling term of the form $g(a+a^{\dagger})(\sigma^{-}+\sigma^{+})$, does not conserve the total number of excitations due to the terms $ga\sigma^{-}$ and $ga^{\dagger}\sigma^{+}$. However, these terms still keep the total parity constant, since they add or remove excitations by an even number. Therefore, most of the aforementioned processes, such as the simultaneous absorption of a single photon by several atoms [171, 174], require the implementation of a generalized Rabi model with the addition of another coupling term of the form $\sigma_{z}(a+a^{\dagger})$, which allows to increase or decrease excitations by an odd number. As a result, one can engineer non-linear processes characterized by the coherent oscillation between an initial state $\left|{i}\right\rangle$ and a final state $\left|{f}\right\rangle$ with a different total number of excitations, such as the simultaneous absorption of a single photon by two atoms, that would be represented by the two states $\left|{i}\right\rangle=\left|{1,g,g}\right\rangle$ and $\left|{f}\right\rangle=\left|{0,e,e}\right\rangle$. This can be achieved if the system parameters are tuned such that $\left|{i}\right\rangle$ and $\left|{f}\right\rangle$ are eigenstates of the bare Hamiltonian with the same energy, allowing one to obtain an effective Hamiltonian that generates the coherent conversion between the two states,
@@
From the point of view of simulations, a recent proposal [186] showed that all the nonlinear processes emerging from the generalized quantum Rabi model — including those breaking parity conservation — can be simulated in JC systems well below the USC regime, using only a single coherent drive. This result is achieved by working in the dressed basis of the atoms driven by the classical field. Let us consider the Hamiltonian of a single driven atom coupled to a cavity mode in the weak-driving regime described by the JC model, and written in the rotating frame of the drive: